A triangular board has a height two thirds as tall as the base. If the area of the board is 75sq cm, what is the length of the base and height?
Answer by mananth(16946)

let the base be x
the height = 2/3 * x = 2x/3
Area of triangle = 1/2 * base * height
1/2 *x*2x/3 = 75
x^2/3 =75
x^2=225
x = sqrt 225
x=15 cm the base
..
height = 2/3 * 15
=10 cm.